OpenStudy (anonymous):

Find the value of $825 invested at 5% with interest compounded twelve times per year for 9 years

OpenStudy (anonymous):

$825.00([1 + 0.05/12]^[12 * 9]) = $825.00([1 + 5/1200]^108) = $825.00([1 + 1/240]^108) = $825.00([241/240]^108) = $825.00(1.566846649) = $1,292.65

OpenStudy (anonymous):

Gee thanks @ihateschool18 Shucks you are welcome

OpenStudy (anonymous):

thanks buddy :)